# Towards Embedding Dynamic Personas in Interactive Robots: Masquerading Animated Social Kinematics (MASK)

### v1

@@ V-B Human Subject Study Results
We examine our hypothesis with classification accuracy among characters based on the results of survey questions S2 and S3. The classification accuracy was reported as $76.7\%$, higher than the threshold set on mentioned hypothesis 2 ($67\%$). The confusion matrix for classification is shown in Figure 3, where we have adjusted the user’s fitting score obtained from the survey question S6. The confusion matrix $T_{ij}$ is defined as $T_{ij}=\frac{\sum_{k}\big{(}s_{f}(k)\cdot\mathds{1}(c_{k}=c_{i})\cdot\mathds{1}(g_{k}=c_{j})\big{)}}{\sum_{k}\big{(}s_{f}(k)\cdot\mathds{1}(g_{k}=c_{j})\big{)}}$, where $k$ is an index of a participants, $s_{f}(k)$ as a fitting score observed from S3, $c_{k}$ as a chosen character from S2, and $g_{k}$ as a persona shown to the participant $k$. We observed that the users could successfully identify the Cowardly Lion from The Wizard of OZ, Minions, and Scrooge characters. However, participants appeared to misclassify the character of Spock, especially between Spock ($0.44$) and Scrooge ($0.39$). Two participants who interacted with persona Spock mentioned that the robot’s firm and blunt behavior with a neutral face leads users to feel that the robot is “grumpy”, making it easily confused with Scrooge. Table II describes the top-3 mentioned keywords for each character from survey question S4. These mentioned keywords from S1 serve as an indicator that users accurately perceive the unique characteristics of persona-driven robots, which also supports our hypothesis.
@@
TABLE II: Top 3 mentioned keywords for each persona from survey question 3.
VI Discussions
In our user studies, we demonstrated that participants could accurately identify and differentiate the robot’s persona across character personas, involving 108 participants. Key findings include discrepancies between user expectations of warm, competent robots and their experiences, particularly with less agreeable or expressive personas, leading to challenges in interaction context comprehension. Additionally, a significant observation was the tendency of participants to anthropomorphize the robot’s actions, actively crafting narratives to understand its behavior, which suggests the importance of user interpretation in interactive robot design. However, our approach faced limitations such as a lack of diverse robotic behaviors in longer interactions, repetitive actions due to a limited number of states, and challenges in group interactions and character representation. These insights point towards the need for a broader observation space, a more complex behavior selection mechanism, and improved design for character portrayal to enhance user engagement and the overall interaction quality with interactive robots. We anticipate that this work will inspire future efforts toward expressive and convincing persona-based robotic agents that provide immersive experiences to people.

### v2

@@ IV-B Human Subject Study Results
We measured the classification accuracy among characters based on user responses to survey questions S5 and S6 to examine how accurately users can identify each persona. We have observed that the participants could successfully identify the correct movie character with an average accuracy of $76.7\%$, $4.6$ times higher than random guess ($16.7\%$). The confusion matrix for classification is shown in Figure 5, where we have adjusted the user’s fitting score obtained from the survey question S6. The confusion matrix $T_{ij}$ is defined as
$T_{ij}=\frac{\sum_{k}\big{(}s_{f}(k)\cdot\mathds{1}(c_{k}=c_{i})\cdot\mathds{1}(g_{k}=c_{j})\big{)}}{\sum_{k}\big{(}s_{f}(k)\cdot\mathds{1}(g_{k}=c_{j})\big{)}}$ (6)
where $k$ is an index of a participants, $s_{f}(k)$ as a fitting score observed from S6, $c_{k}$ as a chosen character from S5, and $g_{k}$ as a persona shown to the participant $k$. We observe that users could successfully identify the Cowardly Lion from The Wizard of OZ, Minions, and Scrooge characters with great accuracy. However, participants appeared to misclassify the character of Spock with Scrooge ($0.39$). Two participants who interacted with the Spock persona mentioned that the robot’s firm and blunt behavior with a neutral face leads users to feel that the robot is “grumpy”, making it easily confused with Scrooge.
Qualitative Result
From S4, the users were asked to mention the keywords associated with each persona before answering S5. Again, the survey was conducted in Korean; the keywords are the closest translations available. Table II describes the top-three mentioned keywords for each character. Participants attributed traits matching each character’s known qualities: “shy” for the Cowardly Lion, “playful” for the Minions, “angry” for Scrooge, and “indifferent” for Spock. These keywords from S4 indicate that users accurately perceive the unique characteristics of persona-driven robots.
Throughout the detailed comments, we observe that most participants could understand the context of the behavior. One participant commented that “Judging by the constant peek-a-boo, [the robot] seemed like a friend with a strong sense of playfulness. Also, since it appeared to yawn when we didn’t move, it seemed to get bored quickly if not engaged, representing a mischief character.” after interacting with the Minion persona. More than half of the participants ($15$ out of $27$) who interacted with Minions persona commented positively about active engagement and playful character. We also observed that some participants created their own stories of the show, inducing a new interpretation of their behaviors. For the cowardly lion persona, one participant commented that “The robot seemed to interpret the action of waving arms as an act of aggression, taking a defensive posture and shedding tears as if it was afraid of humans.”. However, another user reported that “It is impossible to understand the message it is trying to convey, and I am not feeling the interaction.” after interacting with Spock, pointing out the lack of interpretability in unexpressive behaviors.
@@
TABLE II: Top 3 mentioned keywords for each persona from survey question 4 (S4).
V Discussion
Throughout our user studies, we have shown that the participants could successfully recognize and distinguish the robot’s persona in both personalities in Section IV-B (52 participants) and characters in Section IV-B (108 participants). In this section, we discuss the lessons learned during our studies and the main limitations of our work.
Human’s Expectation Towards Robot
We found discrepancies in user expectations versus the robot’s behavior, as most participants were expecting warm and competent robot behaviors. We observed what appears to be a major user expectation mismatch regarding disagreeable personas, resulting in a notable decline in satisfaction levels. This trend indicates an expectation mismatch, as participants were unprepared for disagreeable personas. Furthermore, participants expected a competent robot that could respond to users rapidly and proactively. However, with inexpressive characters, such as the Spock persona where the robot maintained a static state in response to users’ greetings, participants found it challenging to understand the context of the interaction and sometimes believed the robot was malfunctioning. This underlines the need for discreet selection of a robot’s persona or additional cues to inform users in addition to non-verbal behaviors representing personas that consider the user expectations and the interaction context.
Anthropomorphism
We noticed a tendency among participants to anthropomorphize the robot’s actions as participants crafted their own narratives and interpretations of the robot’s behaviors within the presented scenarios. For instance, one participant interpreted that the robot perceived the user’s hand waving as a threat after observing defensive motions from the robot. Additionally, two participants noted that the robot with a playful character (Minion) seemed to exhibit boredom when not greeted. We believe the users had an engaging interaction by actively interpreting the robot’s behavior and creating their own story in the show. This suggests that our research offers insights into how anthropomorphism influences user engagement and perception, highlighting the significant role of user interpretation in designing interactive robots.
Limitations
Despite the proposed method’s capability of mimicking characters or embodying personalities, our system currently suffers from some limitations and shortcomings. Most significantly, users reported a lack of diversity in robotic behavior for interactions that lasted beyond a minute. The proposed system does not offer a sufficient variety of actions and observations. Furthermore, the limited number of states available for each persona has resulted in users perceiving the robot as engaging in repetitive behavior. This highlights a significant drawback to our system which should be considered in future iterations: added novelty and robotic behaviors for longer interactions will require a larger observation space to detect more non-verbal cues, and a library of robot actions that can be drawn by a more complex state machine within the behavior selection engine. Ultimately, a state machine architecture is not preferred as its complexity increases exponentially with added states - a new method of selecting appropriate behaviors would need to be implemented.
In addition, in group settings, we observed a delay in responding to the user’s actions; when multiple people attempted to interact with the robot simultaneously, the robot could not interact with all of the users simultaneously. We have seen cases where users’ curiosity scores decreased as the robot was still in states that responded to other participants. A more complex turn-taking system [23] should be considered rather than a single curiosity score for enhancing the interaction quality. Finally, in character-based persona interaction, motions are not tailored to reflect the character, and there is a difference in the appearance of the character and the robot (e.g., AMBIDEX [14] is too big to look like Minion). These two factors contributed to participants noticing a gap between the intended character portrayal and the robot. Therefore, a review of the method for stylizing movements and potential design modifications should be taken into account.
VI Conclusion
We have presented a system which we call MASK for interacting with humans non-verbally based on personas that can take on the likeness of personalities or film characters. This system integrates several components (perception engine with body pose estimation, behavior selection engine that incorporates an LLM for behavior selection, and action library with actions created through a motion retargeting pipeline) to take steps towards generating autonomous, interactive robot characters. As evidenced through user subject studies, we saw that conditioning robotic behavior with a state, an observation, and a persona can enable users to experience a convincing persona in interactive agents. We observed that users could recognize the intended character in both personality- and film-character-based personas. Despite the technical limitations of our system, such as lack of diversity in behavior and delay of response in group setting scenarios, we have shown preliminary steps towards bringing autonomous character-like agents into the physical world. We anticipate that this work will inspire future efforts toward expressive and convincing persona-based robotic agents that provide captivating experiences to people.